# Thermodynamic Cost of Random-Time Protocols

### v2

@@ Abstract
Systems that are driven by a randomly timed, external protocol can seemingly violate the second law of thermodynamics. We show that this thermodynamic paradox is resolved if the outcome of the random time is stored in a memory device. Specifically, we show that the average work required to erase the memory is always larger than the average work gained from the protocol. We also discuss experimentally implementable setups that measure random times directly without continuous monitoring. Taken together, this paper discusses the relationship between temporal information and thermodynamics. This framework is relevant for external protocols employing random times, such as, stochastic resetting protocols and cyclically driven heat engines that use randomly timed protocols.
@@ I Introduction
Furthermore we introduce in this paper a non-autonomous demon that directly measures first-passage times. Although first-passage demons have been considered before in the literature, see Refs. [7, 8, 9, 10, 11], these non-autonomous demons in fact measure the whole trajectory of the system up to the first-passage time through continuous monitoring. In setups of non-autonomous demons, continuous monitoring requires a very large number of measurements, as continuous monitoring is implemented through a protocol that measures the state of the system at regularly spaced time intervals. Here, instead, we develop a simpler setup for which a non-autonomous demon directly measures the first-passage time using a single measurement. We present concrete examples of experimentally implementable setups that measure first-passage times directly without continuous monitoring.
The paper is structured as follows. In Sec. II, we introduce the system setup for externally controlled protocols that operate with random times, focusing on the isothermal case. In Sec. III, we argue, by analogy with Szilard’s engine, that a non-autonomous demon implementing a random-time protocol requires a memory device to run the protocol. In Sec. IV, we show that the storage of a random time is sufficient to derive the second law of thermodynamics for systems controlled by random-time protocols, further confirming that a memory device is needed in these setups. The Secs. V to VII explore a paradigmatic example of a temporal information engine, which is the random-time version of Szilard’s engine as illustrated in Fig. 1. This model is arguably the simplest example of a temporal information engine. It is a simplification of previous setups based on continuous Maxwell demons [7, 8, 9, 10, 11], as the demon measures the random time directly without continuous monitoring. In Sec. V, we consider the random-time Szilard engine in discrete time and without measurement errors, and in Sec. VI we consider the corresponding model in continuous time. In Sec. VII we develop a simple model for random-time detection without continuously monitoring the system, confirming that the model is physically sound. In Sec. VIII we show that stochastic heat engines that use random-time protocols satisfy the Carnot bound on their efficiency, which is further evidence in support of the claim that memory storage is necessary in random-time protocols. Section IX compares the setups of Maxwell demons that continuously monitor a system with those that directly measure first-passage times. We end the paper with a discussion in Sec. X and a few appendices.
@@ IV Second law of thermodynamics for non-autonomous demons that use random times
We note that the Eq. (7) is a stochastic thermodynamics result. It can be derived in the setup of Markov processes that satisfy local detailed balance (see Appendix A). An alternative derivation of the inequality (7) that is based on continuously monitoring the system is presented in Ref. [10]. Following Refs. [41, 11], the inequality (7) can be further refined by taking into account the unavailable information. However, our purpose is to derive the second law of thermodynamics for randomly timed protocols, and for this purpose the inequality (7) suffices.
@@ V Random-time Szilard engine: discrete version
Note that sleeping demons simplify setups based on continuous Maxwell demons. Continuous Maxwell demons measure the full trajectory $X^{T}_{0}$ through continuous monitoring [7, 8, 9, 10]. Instead, the sleeping demon directly measures $T$ without continuous monitoring. This is possible within classical physics as we show in Sec. LABEL:sec7.
@@ VII Random-time detectors
The random-time Szilard engine of Fig. 1 contains a random-time detector, which so far we have assumed to be perfect. Instead, we develop now a physical model for the detector in the discrete random-time Szilard engine.
@@
Note that the function of the detector in both of the above examples is to awaken the demon. Once the demon is alerted, it measures the time $T$, stores $T$ in a memory device, and initiates the protocol $\lambda_{T}(s)$. One caveat here is a that the demon needs a clock to measure time. We elaborate on this in the discussion section of the paper.
@@ IX.1 Continuous monitoring
Let us first revisit the continuous monitoring formalism, see Refs. [7, 8, 9, 10, 11]. In the continuous monitoring framework, the demon measures $X$ at regularly spaced time intervals $0,\delta t,2\delta t,\ldots$. The demon inserts a moveable wall at position $\ell-\tilde{\ell}$ the first time when $X(n\delta t)\in[\ell-\tilde{\ell}$$\ell$] (notice that it is more efficient to measure a binary variable that indicates whether $X(n\delta t)\in[\ell-\tilde{\ell}$$\ell$]). To avoid errors, we need to consider the limit of $\delta t\rightarrow 0$ which is the limit of continuous monitoring.
@@ X Discussion
To illustrate the consequences of temporal information on nonequilibrium thermodynamics, we have introduced a non-autonomous Maxwell demon that directly measures the random time $T$, and which we called the sleeping demon. We have developed experimentally implementable setups that measure first-passage times directly, unlike the conventional approach based on continuous monitoring that is typically employed in the literature [39, 45, 10, 7, 8, 9, 11]. As in these setups there is a single measurement, the cost of measurement is fixed and independent on the size of the system. This is different from setups based on continuous measurements, for which multiple measurements are performed, and therefore the cost of measurement grows with the system size.
@@
Szilard engines have been implemented in electronic devices, such as single electron boxes [51], which raises the question on how to describe temporal information in quantum mechanics. We expect conceptual difficulties due to the quantum Zeno effect [52, 53, 54, 55]. There is also a possibility to define first-passage times in Bohmian mechanics [56, 57], which may yield an alternative perspective on temporal information in quantum mechanics.
@@ Acknowledgements.
Note that it is possible to refine the result (31) into an equality, see the approach in Ref. [41]. However, for our purpose of deriving the second law for randomly timed protocols the result (31) suffices.
We derive the bound (31) for Markov chains in discrete time, even though the results also apply in continuous time. Our derivations are inspired by Refs. [38, 39, 40], albeit we extend their arguments to protocols that rely on temporal information. For the analysis below, we require that $T$ is a stopping time (i.e., $T$ is independent of the values of $X$ after $T$).

### v3

@@ Abstract
Systems that are driven by a randomly timed, external protocol can seemingly violate the second law of thermodynamics. We show that this thermodynamic paradox is resolved if the outcome of the random time is stored in a memory device. Specifically, we show that the average work required to erase the memory is always larger than the average work gained from the protocol. We also discuss concrete setups that measure random times directly without continuous monitoring. Taken together, this paper discusses the relationship between temporal information and thermodynamics. This framework is relevant for external protocols employing random times, such as, stochastic resetting protocols and cyclically driven heat engines that use randomly timed protocols.
@@ I Introduction
Furthermore we introduce in this paper a non-autonomous demon that directly measures first-passage times. Although first-passage demons have been considered before in the literature, see Refs. [7, 8, 9, 10, 11], these non-autonomous demons in fact measure the whole trajectory of the system up to the first-passage time through continuous monitoring. In setups of non-autonomous demons, continuous monitoring requires a large number of measurements, as continuous monitoring is implemented through a protocol that measures the state of the system at regularly spaced time intervals. Here, instead, we develop a simpler setup for which a non-autonomous demon directly measures the first-passage time using a single measurement. We present concrete examples of setups that measure first-passage times directly without continuous monitoring.
The paper is structured as follows. In Sec. II, we introduce the system setup for externally controlled protocols that operate with random times, focusing on the isothermal case. In Sec. III, we argue, by analogy with Szilard’s engine, that a non-autonomous demon implementing a random-time protocol requires a memory device to run the protocol. In Sec. IV, we show that the storage of a random time is sufficient to derive the second law of thermodynamics for systems controlled by random-time protocols, further confirming that a memory device is needed in these setups. As we demonstrate, the second law of thermodynamics follows from combining Landauer’s erasure principle [42, 27, 28] with the formalism of information thermodynamics for non-autonomous Maxwell demons with a path-dependent protocol [38, 39, 40]. The Secs. V to VII explore a paradigmatic example of a temporal information engine, which is the random-time version of Szilard’s engine as illustrated in Fig. 1. This model is arguably the simplest example of a temporal information engine. It is a simplification of previous setups based on continuous Maxwell demons [7, 8, 9, 10, 11], as the demon measures the random time directly without continuous monitoring. In Sec. V, we consider the random-time Szilard engine in discrete time and without measurement errors, and in Sec. VI we consider the corresponding model in continuous time. In Sec. VII we develop a simple model for random-time detection without continuously monitoring the system, confirming that the model is physically sound. In Sec. VIII we show that stochastic heat engines that use random-time protocols satisfy the Carnot bound on their efficiency, which is further evidence in support of the claim that memory storage is necessary in random-time protocols. Section IX compares the setups of Maxwell demons that continuously monitor a system with those that directly measure first-passage times. We end the paper with a discussion in Sec. X and a few appendices.
@@ IV Second law of thermodynamics for non-autonomous demons that use random times
We note that the Eq. (7) is a stochastic thermodynamics result. It can be derived in the setup of Markov processes that satisfy local detailed balance (see Appendix A). An alternative derivation of the inequality (7) that is based on continuously monitoring the system is presented in Ref. [10]. This is because at the level of mathematical formulation, both the sleeping demon and the conventional continuous demon can be described in a unified setup of error-free non-autonomous Maxwell demon with a path-dependent protocol $\lambda(t;X^{t}_{0})$ [38, 39, 40]. Furthermore, following Ref. [41], the inequality (7) can be further refined by taking into account the unavailable information. However, our purpose is to derive the second law of thermodynamics for randomly timed protocols, and for this purpose the inequality (7) provides the most parsimoniuous explanation.
@@ V Random-time Szilard engine: discrete version
Note that sleeping demons simplify setups based on continuous Maxwell demons. Continuous Maxwell demons measure the full trajectory $X^{T}_{0}$ through continuous monitoring [7, 8, 9, 10]. Instead, the sleeping demon directly measures $T$ without continuous monitoring. This is possible within classical physics as we show in Sec. VII.
@@ VII Random-time detectors
The random-time Szilard engine of Fig. 1 contains a random-time detector, which so far we have assumed to be perfect. We develop now physical models for random-time detectors. We first consider the case of a discrete random-time Szilard engine.
@@
Note that the function of the detector in both of the above examples is to awaken the demon. Once the demon is alerted, it measures the time $T$, stores $T$ in a memory device, and initiates the protocol $\lambda_{T}(s)$. One caveat here is that the demon needs a clock to measure time. We elaborate on this in the discussion section of the paper.
@@ IX.1 Continuous monitoring
Let us first revisit the continuous monitoring formalism, see Refs. [7, 8, 9, 10, 11] and  [39, 41, 45]. In the continuous monitoring framework, the demon measures $X$ at regularly spaced time intervals $0,\delta t,2\delta t,\ldots$. The demon inserts a moveable wall at position $\ell-\tilde{\ell}$ the first time when $X(n\delta t)\in[\ell-\tilde{\ell}$$\ell$] (notice that it is more efficient to measure a binary variable that indicates whether $X(n\delta t)\in[\ell-\tilde{\ell}$$\ell$]). To avoid errors, we need to consider the limit of $\delta t\rightarrow 0$ which is the limit of continuous monitoring.
@@ X Discussion
To illustrate the consequences of temporal information on nonequilibrium thermodynamics, we have introduced a non-autonomous Maxwell demon that directly measures the random time $T$, and which we called the sleeping demon. We have developed setups that measure first-passage times directly, unlike the conventional approach based on continuous monitoring that is typically employed in the literature [10, 7, 8, 9, 11]. In the direct approach, the cost of measurement is fixed and independent on the size of the system. This is different from setups based on continuous measurements, for which multiple measurements are performed, and therefore the cost of measurement grows with the system size.
@@
Szilard engines have been implemented in electronic devices, such as single electron boxes [51], which raises the question on how to describe temporal information in quantum mechanics. We expect some complications due to the quantum Zeno effect [52, 53, 54, 55]. There is also a possibility to define first-passage times in Bohmian mechanics [56, 57], which may yield an alternative perspective on temporal information in quantum mechanics.
@@ Acknowledgements.
Note that it is possible to refine the result (31) into an equality following the approach in Ref. [41]. However, for our purpose of deriving the second law for randomly timed protocols the result (31) suffices.
We derive the bound (31) for Markov chains in discrete time, even though the results also apply in continuous time. Our derivations are based on the framework of error-free non-autonomous Maxwell demons with a path-dependent protocol $\lambda(t;X^{t}_{0})$, see Refs. [38, 39, 40]. For the analysis below, we require that $T$ is a stopping time (i.e., $T$ is independent of the values of $X$ after $T$).